OpenStudy (anonymous):

sinx/1-cosx+sinx/1+cosx=2cscx

OpenStudy (mertsj):

What are you supposed to do with that?

OpenStudy (anonymous):

Prove it

OpenStudy (mertsj):

\[\frac{\sin x}{1-\cos x}+\frac{\sin x}{1+\cos x}=2\csc x\]

OpenStudy (mertsj):

\[\frac{\sin x(1+\cos x)}{(1-\cos x)(1+\cos x)}+\frac{\sin x(1-\cos x)}{(1+\cos x)(1-\cos x)}\]

OpenStudy (mertsj):

\[\frac{\sin x(1+\cos x)}{1-\cos ^2x}+\frac{\sin x(1-\cos x)}{1-\cos ^2x}\]

OpenStudy (mertsj):

\[\frac{\sin x(1+\cos x)}{\sin ^2x}+\frac{\sin x(1-\cos x)}{\sin ^2x}\]

OpenStudy (mertsj):

\[\frac{1+\cos x}{\sin x}+\frac{1-\cos x}{\sin x}=\frac{2}{\sin x}=2\csc x\]